# The Impact of AI Innovations on U.S. Occupations

### 07 Dec 2023 v1

@@ Abstract
AI’s impact has traditionally been assessed in terms of occupations. However, an occupation is comprised of interconnected tasks, and it is these tasks, not occupations themselves, that are affected by AI. To evaluate how tasks may be impacted, previous approaches utilized subjective manual annotations or coarse-grained matching with patents. Leveraging recent advancements in machine learning, we replace coarse-grained matching with more precise deep learning approaches. Introducing the AI Impact (AII) measure, we employ Deep Learning Natural Language Processing to automatically identify AI patents that impact various occupational tasks at scale. Our methodology relies on a comprehensive dataset of 19,498 task descriptions and quantifies AI’s impact through analysis of 12,984 AI patents filed with the United States Patent and Trademark Office (USPTO) between 2015 and 2020. Our observations reveal that the impact of AI on occupations defies simplistic categorizations based on task complexity, challenging the conventional belief that the dichotomy between basic and advanced skills alone explains the effects of AI. Instead, the impact is intricately linked to specific skills, whether basic or advanced, associated with particular tasks. For instance, while basic skills like scanning items may be affected, others like cooking may not. Similarly, certain advanced skills, such as image analysis in radiology, may face impact, while skills involving interpersonal relationships may remain unaffected. Furthermore, the influence of AI extends beyond knowledge-centric regions. Regions in the U.S. that heavily rely on industries susceptible to AI changes, often characterized by economic inequality or a lack of economic diversification, will experience notable AI impact.
@@ 1 Introduction
The second class measures AI impact on tasks rather than occupations. This concept is illustrated by Brynjolfsson et al.’s Suitability for Machine Learning (SML) method [3], which measures the impact of AI using a comprehensive set of 18,156 tasks spanning 964 different occupations. However, SML relies on the assessments of crowdworkers to determine the suitability of specific tasks for machine learning. This reliance may introduce subjective biases from annotators (e.g., varying levels of expertise or cultural factors may lead to inconsistencies), and poses challenges in terms of scalability. Also, similar to Frey and Osborne’s and AIOE, SML is limited by the static nature of its one-time manual labeling. As technology advances and new capabilities emerge such as Large Language Models (LLMs), relying solely on a fixed set of abilities or subjective assessments of tasks suitability for automation becomes increasingly inadequate. For example, a copywriter is likely to be impacted by LLMs [16]. However, if one were to examine copywriters at different points in time, such as in 2010 or 2015, the impact of AI on them would not be constant but would drastically change since language models were not as powerful back in 2010 as they are today. To fully capture the impact of a fast moving technology such as AI, therefore, it is crucial for methods to be adaptable to the ever-evolving technological advancements. To gauge the likelihood of automation, it is essential to identify which systems are poised for construction and commercialization. Patents are a typical source in scholarly work to identify emerging technological innovations [17, 18, 19]. They provide insights into emerging systems and technologies [20], leading Webb to study AI innovations by comparing occupational task descriptions with patent titles [2]. This method, employing a dictionary approach to identify verb-noun pairs associated with both tasks and patent titles, has limitations though. Another method, similar in its approach to Webb’s term matching, employs a normalized term matching approach to determine the similarity between tasks and patents [21], and does so to the specific area of robotics rather than AI. The problem is that term matching approaches do not account for the semantic meaning of words or their order [22] and, as such, they either miss relevant patents or return spurious task-patent matches, as detailed in Tables C.1 and C.2.
@@
To overcome these limitations, we introduced and validated the AI Impact (AII) measure. AII utilizes 19,498 task descriptions and assesses AI’s impact through innovations found in 12,984 AI patents filed with the United States Patent and Trademark Office (USPTO) from 2015 to 2020. For consistency with prior methods, our analysis extends up to 2020 (additional analyses up to 2022 are in §D). Built on Sentence-T5 (ST5)[23], a natural language processing-focused deep learning framework (Figure 6), the method gauges semantic similarity between occupation task descriptions and patent descriptions (explained in §4.1). The AII score is calculated in three steps. Firstly, the method identifies the most similar patent for each task based on maximum cosine similarity. Secondly, it categorizes a task as AI-impacted if its similarity with the most similar patent surpasses a threshold at the 90th percentile, as previous literature suggested [20] and this work further empirically validated (§B). Given that patents can be general, our approach addresses this by using the text not only in titles, as previous approaches did [2], but also in abstracts, which are more likely to contain application domains or references to specific tasks (illustrated in Table B.1). Finally, the AII score for an occupation is computed by dividing the number of tasks impacted by AI patents by the total tasks for that occupation. For insights into AI’s economic and spatial ramifications, the method aggregates AII scores for each occupation at industry sector- or regional-levels (elaborated in §4.2).
@@ 1 Cardiovascular Technologists and Technicians Pile-Driver Operators
Least-impacted Occupations. The least-impacted occupations mainly consist of blue-collar occupations such as pile-driver operators, agricultural graders and sorters, floor sanders, aircraft cargo handling supervisors, and dredge operators Indeed, the AII score, binned by education levels, shows that the lowest impact is seen for jobs requiring a high school diploma or less (Figure 1). These occupations are primarily found in agriculture, transportation, accommodation and food services, and construction sectors, where the core tasks and responsibilities revolve around physical and manual labor and typically do not require a wide range of complex mental or physical activities, nor do they involve abstract reasoning. In agriculture, least-impacted tasks involve food inspection and dairy management for livestock. In transportation, least-impacted tasks involve scheduling and resource allocation for airline operations, engine sound control, and vehicle dispatching. In accommodation and food services, tasks include monitoring and recording food temperatures. In construction, least-impacted tasks involve the maintenance and operation of equipment and machinery. In addition to these sectors, by examining occupations with nearly zero AII scores, we found another least impacted set of occupations. Contrary to the previous least impacted occupations which involve manual labor or dexterity, managerial occupations typically require human interactions, having tasks ranging from contract negotiation to proposal review to internal assessments and audits.
@@ 2.2 Most- and Least-Impacted Industry Sectors
Least-impacted sectors, on the other hand, include construction, accommodation and food services, real estate, education, public administration, and managerial positions across industry sectors. These sectors have experienced a significantly low rate of impact. From the thematic analysis, three possible explanations emerge.
The first explanation centers on the fact that occupations within these sectors are frequently associated with low-skilled and physically demanding labor. For example, in construction, tasks can span from assembling solar panels to maintaining pipe systems to operating a variety of drills, all of which require physical labor.
The second explanation is that occupations within these sectors require basic and non-specialized skills. For example, in accommodation services, the tasks of waiters and baristas generally involve minimal specialized knowledge or vocational training, such as serving food and drinks. Likewise, in the real estate, brokers or clerks primarily require training in overseeing transactions and tasks related to office operations.
The third explanation is that occupations within these sectors often involve a high level of skill and require extensive interpersonal interactions. For example, in managerial positions (e.g., CEOs), tasks typically include responsibilities such as delegating tasks, attending events and meetings, and negotiating contracts, all of which heavily rely on human interpersonal communication. Similarly, in education, a teacher’s role primarily centers on delivering educational materials, a task that requires a higher level of education and specialized knowledge. Finally, in the sector of public administration, certain legal occupations involve tasks that are client-facing or require the design of bespoke legal solutions, making them the least-impacted occupations, a finding in line with previous qualitative analysis on legal occupations [28].
@@ 3.3 Limitations and Future Work
This work has four primary limitations. First, our method relies on patent abstracts to provide a finer-grained understanding of occupational tasks, mainly from the US-oriented USPTO dataset. Future research should explore potential cultural differences in how patents are written and used in other contexts, such as the European Union.
Second, it is important to acknowledge that the existing portfolio of USPTO patents may not comprehensively cover all tasks associated with a particular occupation. To address this limitation, one potential approach could involve supplementing patent data with other sources, including research papers and code repositories [48].
Third, our assumption is that the likelihood of an AI system being built is determined by whether it is patented. While generally true, there are exceptions. A patented system may not be built, as the patent could be intended for defensive or offensive purposes [49, 50]. Conversely, a non-patented system might still be constructed, with its design protected by secrecy or trademarks [51, 52].

### 15 Dec 2023 v2

@@ Abstract
AI’s impact has traditionally been assessed in terms of occupations. However, an occupation is comprised of interconnected tasks, and it is these tasks, not occupations themselves, that are affected by AI. To evaluate how tasks may be impacted, previous approaches utilized manual annotations or coarse-grained matching. Leveraging recent advancements in machine learning, we replace coarse-grained matching with more precise deep learning approaches. Introducing the AI Impact (AII) measure, we employ Deep Learning Natural Language Processing to automatically identify AI patents that impact various occupational tasks at scale. Our methodology relies on a comprehensive dataset of 19,498 task descriptions and quantifies AI’s impact through analysis of 12,984 AI patents filed with the United States Patent and Trademark Office (USPTO) between 2015 and 2020. Our observations reveal that the impact of AI on occupations defies simplistic categorizations based on task complexity, challenging the conventional belief that the dichotomy between basic and advanced skills alone explains the effects of AI. Instead, the impact is intricately linked to specific skills, whether basic or advanced, associated with particular tasks. For instance, while basic skills like scanning items may be affected, others like cooking may not. Similarly, certain advanced skills, such as image analysis in radiology, may face impact, while skills involving interpersonal relationships may remain unaffected. Furthermore, the influence of AI extends beyond knowledge-centric regions. Regions in the U.S. that heavily rely on industries susceptible to AI changes, often characterized by economic inequality or a lack of economic diversification, will experience notable AI impact.
@@ 1 Introduction
The second class measures AI impact on tasks rather than occupations. This concept is illustrated by Brynjolfsson et al.’s Suitability for Machine Learning (SML) method [3], which measures the impact of AI using a comprehensive set of 18,156 tasks spanning 964 different occupations. However, SML relies on the assessments of crowdworkers to determine the suitability of specific tasks for machine learning. This reliance may introduce subjective biases from annotators (e.g., varying levels of expertise or cultural factors may lead to inconsistencies), and poses challenges in terms of scalability. Also, similar to Frey and Osborne’s and AIOE, SML is limited by the static nature of its one-time manual labeling. As technology advances and new capabilities emerge such as Large Language Models (LLMs), relying solely on a fixed set of abilities or subjective assessments of tasks suitability for automation becomes increasingly inadequate. For example, a copywriter is likely to be impacted by LLMs [16]. However, if one were to examine copywriters at different points in time, such as in 2010 or 2015, the impact of AI on them would not be constant but would drastically change since language models were not as powerful back in 2010 as they are today. To fully capture the impact of a fast moving technology such as AI, therefore, it is crucial for methods to be adaptable to the ever-evolving technological advancements. To gauge the likelihood of automation, it is essential to identify which systems are poised for construction and commercialization. Annual business surveys serve as a source for measuring the adoption of automation. However, they may be subject to biases (e.g., respondents may over-report positive aspects, prioritize certain business operations, potentially neglecting others, interpret survey questions subjectively without standardized criteria) and are infrequently updated [17]. An alternative, more objective source is patents. Patents are a typical source in scholarly work to identify emerging technological innovations [18, 19, 20]. They provide insights into emerging systems and technologies [21], leading Webb to study AI innovations by comparing occupational task descriptions with patent titles [2]. This method, employing a dictionary approach to identify verb-noun pairs associated with both tasks and patent titles, has limitations though. Another method, similar in its approach to Webb’s term matching, employs a normalized term matching approach to determine the similarity between tasks and patents [22], and does so to the specific area of robotics rather than AI. The method most similar to ours was proposed in [23, 24], in which, using word embedding, patents are matched with broad occupation categories from the American Community Survey. However, that level of categorization is not suitable for researching the characteristics of specific jobs. Overall, this second class of approaches has used either term matching or word embedding. The problem is that term matching does not capture the semantic meaning of words (e.g., it does not distinguish between ‘bank’ as in ‘data bank’ or as in a financial institution) [25], and word embedding does not account for word ordering (e.g., ‘data entry and analysis’ and ‘analysis of entry data’ are considered similar based on word embeddings yet are two different tasks). As a result, these methods either miss relevant patents or return spurious task-patent matches, as detailed in Tables C.1 and C.2.
@@
To overcome these limitations, we introduced and validated the AI Impact (AII) measure. AII utilizes 19,498 task descriptions from O*NET and assesses AI’s impact through innovations found in 12,984 AI patents filed with the United States Patent and Trademark Office (USPTO) from 2015 to 2020. For consistency with prior methods, our analysis extends up to 2020 (additional analyses up to 2022 are in §D). Built on Sentence-T5 (ST5)[26], a natural language processing-focused deep learning framework (Figure 6), the method gauges semantic similarity between occupation task descriptions and patent descriptions (explained in §4.1) by embedding not individual words but the entire document (e.g., the entire patent’s abstract), allowing for considering both semantic meaning and word ordering. The AII score is calculated in three steps. Firstly, the method identifies the most similar patent for each task based on maximum cosine similarity. Secondly, it categorizes a task as AI-impacted if its similarity with the most similar patent surpasses a threshold at the 90th percentile, as previous literature suggested [21] and this work further empirically validated (§B). We select, for each task, the closest patent rather than counting the number of closely linked patents. This approach provides a more targeted understanding of the specific innovations or solutions directly relevant to that particular task, avoiding dilution of the analysis with potentially less pertinent or peripheral patents. Also, given that patents can be general, our approach addresses this by using the text not only in titles, as previous approaches did [2], but also in abstracts, which are more likely to contain application domains or references to specific tasks (illustrated in Table B.1). Finally, the AII score for an occupation is computed by dividing the number of tasks impacted by AI patents by the total tasks for that occupation. For insights into AI’s economic and spatial ramifications, the method aggregates AII scores for each occupation at industry sector- or regional-levels (elaborated in §4.2).
@@ 1 Cardiovascular Technologists and Technicians Pile-Driver Operators
Least-impacted Occupations. The least-impacted occupations mainly consist of blue-collar occupations such as pile-driver operators, agricultural graders and sorters, floor sanders, aircraft cargo handling supervisors, and dredge operators. Indeed, the AII score, binned by education levels, shows that the lowest impact is seen for jobs requiring a high school diploma or less (Figure 1). These occupations are primarily found in agriculture, transportation, accommodation and food services, and construction sectors, where the core tasks and responsibilities revolve around physical and manual labor and typically do not require a wide range of complex mental or physical activities, nor do they involve abstract reasoning. In agriculture, least-impacted tasks involve food inspection and dairy management for livestock. In transportation, least-impacted tasks involve scheduling and resource allocation for airline operations, engine sound control, and vehicle dispatching. In accommodation and food services, tasks include monitoring and recording food temperatures. In construction, least-impacted tasks involve the maintenance and operation of equipment and machinery. In addition to these sectors, by examining occupations with nearly zero AII scores, we found another least impacted set of occupations. Contrary to the previous least impacted occupations which involve manual labor or dexterity, managerial occupations typically require human interactions, having tasks ranging from contract negotiation to proposal review to internal assessments and audits.
@@ 2.2 Most- and Least-Impacted Industry Sectors
Least-impacted sectors, on the other hand, include construction, accommodation and food services, real estate, education, public administration, and finance and insurance. These sectors have experienced a significantly low rate of impact. From the thematic analysis, three possible explanations emerge.
The first explanation centers on the fact that occupations within certain sectors are often associated with low-skilled or physical/manual labor. For instance, in construction, tasks range from assembling solar panels to maintaining pipe systems to operating various drills, all of which require physical labor. Additionally, manual dexterity is challenging to automate. In public administration, numerous occupations, such as police officers and firefighters, still require manual skills.
The second explanation is that occupations within certain sectors demand basic and non-specialized skills. For instance, in accommodation services, the tasks of waiters and baristas typically involve minimal specialized knowledge or vocational training, such as serving food and drinks. Similarly, in real estate, brokers or clerks primarily require training in overseeing transactions and handling tasks related to office operations.
The third explanation is that occupations within certain sectors often require a high level of skill and involve extensive interpersonal interactions. For instance, in managerial positions (e.g., CEOs), tasks typically encompass responsibilities such as delegating tasks, attending events and meetings, and negotiating contracts - all of which heavily rely on human interpersonal communication. Similarly, in education, a teacher’s role primarily revolves around delivering educational materials in person, a task that demands both physical presence and a higher level of education and specialized knowledge. In the legal sector, impacted occupations include those involved in drafting legal documents or transcribing pretrial and trial proceedings, such as court reporters, in alignment with previous qualitative analyses on legal occupations [31]. Conversely, roles requiring the design of bespoke legal solutions remain unaffected. Notably, client-facing tasks in the legal sector are also not impacted, resulting in an overall expectation of limited impact on the legal sector. Finally, in the financial sector, there are still occupations that necessitate human interactions, such as clerks, sales agents, and tellers.
@@ 3.3 Limitations and Future Work
This work has six primary limitations. First, our measure reflects whether a patent impacts a task but cannot distinguish whether the impact means replacing the task or augmenting it, which has been a long standing debate in the literature [23, 51, 4, 52, 53].
Second, our analysis is based on currently existing tasks, capturing their potential automation, and cannot be used to study the emergence of new tasks as a result of that automation (referred to in previous work as ‘work augmentation’ [23]).
Third, our method relies on patent abstracts to provide a finer-grained understanding of occupational tasks, mainly from the US-focused USPTO dataset. Future research should explore potential cultural differences in how patents are written and used in other contexts, such as the European Union.
Fourth, it is important to acknowledge that the existing portfolio of USPTO patents may not comprehensively cover all the innovations that may impact a particular occupation. To address this limitation, one potential approach could involve supplementing patent data with other sources, including research papers and code repositories [54].
Fifth, our assumption is that the likelihood of an AI system being built is determined by whether it is patented. While generally true, there are exceptions. A patented system may not be built, as the patent could be intended for defensive or offensive purposes [55, 56]. Conversely, a non-patented system might still be constructed, with its design protected by secrecy or trademarks [57, 58]. Furthermore, patenting rates vary among sectors [59], and there is a lag between an innovation being patented and its use and impacts diffusing across the economy [54].